# Displacement-Squeeze receiver for BPSK displaced squeezed vacuum states surpassing the coherent-states Helstrom bound under imperfect conditions

### v1

@@ Appendix A The Wigner Function of Displaced Squeezed Vacuum States
The Displaced Squeezed States (DSSs) discussed in this paper belong to single-mode Gaussian states, whose Wigner function is a bivariate Gaussian distribution fully characterized by two parameters: the displacement vector $\vec{d}$ and the covariance matrix $\mathbf{V}$. Given a DSS denoted as $\left|\psi\right\rangle$, its quantum mechanical expression is
@@
$\mathbf{V}$ $=\begin{bmatrix}V_{xx}&V_{xp}\\ V_{px}&V_{pp}\end{bmatrix}$ (42)
$=\begin{bmatrix}\left\langle(\Delta\hat{X})^{2}\right\rangle&\frac{1}{2}\left\langle\left\{\Delta\hat{X},\Delta\hat{P}\right\}\right\rangle\\[4.0pt] \frac{1}{2}\left\langle\left\{\Delta\hat{X},\Delta\hat{P}\right\}\right\rangle&\left\langle(\Delta\hat{P})^{2}\right\rangle\end{bmatrix}$,
@@ Appendix B The Helstrom bound and SQL of S-BPSK under phase diffusion condition
$P_{\text{SQL}}^{\text{DSS}}\left(N,\sigma\right)$ $=\int_{0}^{+\infty}\mathrm{d}x\cdot P\left(x|\rho_{i}^{\text{pd}}\right)$ (45)
@@
$=\int_{\mathbf{R}}\mathrm{d}\phi\cdot g_{\sigma}(\phi)\cdot P_{\text{SQL}}^{\text{DSS}}(\phi)$
@@
$P_{\text{SQL}}^{\text{DSS}}(\phi)=\int_{0}^{+\infty}\mathrm{d}x\cdot P\left(x|\rho_{i,\phi}\right)=\int_{0}^{+\infty}\mathrm{d}x\int_{\mathbf{R}}\mathrm{d}p\cdot W\left(x,p|\rho_{i,\phi}\right)$ (46)
@@
$P_{\text{HB}}^{\text{DSS}}\left(N,\sigma\right)=\frac{1}{2}\left(1-\left\|P_{0}\rho_{0}^{\text{pd}}-P_{1}\rho_{1}^{\text{pd}}\right\|_{1}\right)$ (47)
where $\left\|A\right\|_{1}=\mathbf{Tr}\left\{A^{\dagger}A\right\}$ denotes the trace norm, and $P_{0},P_{1}$ denote prior probabilities. This expression applies to the calculation of the Helstrom bound for binary states, whether pure or mixed.
@@ Acknowledgements.
This work was supported by Independent Innovation Science Fund Program of National University of Defense Technology, China (Grant No. 22-ZZCX-036), Key research&development program of Guangxi (Grant No. GuiKeAB23075155), the National Science Foundation for Young Scientists of China (Grant No. 12204539), and the Key Research & Development Program of Guangxi (Grant No. AB23075112).

### v2

@@ Appendix A The Wigner Function of Displaced Squeezed Vacuum States
The displaced squeezed vacuum states discussed in this paper belong to single-mode Gaussian states, whose Wigner function is a bivariate Gaussian distribution fully characterized by two parameters: the displacement vector $\vec{d}$ and the covariance matrix $\mathbf{V}$. Given a displaced squeezed vacuum states denoted as $\left|\psi\right\rangle$, its quantum mechanical expression is
@@
$\mathbf{V}=\begin{bmatrix}V_{xx}&V_{xp}\\ V_{px}&V_{pp}\end{bmatrix}\\ =\begin{bmatrix}\left\langle(\Delta\hat{X})^{2}\right\rangle&\frac{1}{2}\left\langle\left\{\Delta\hat{X},\Delta\hat{P}\right\}\right\rangle\\[4.0pt] \frac{1}{2}\left\langle\left\{\Delta\hat{X},\Delta\hat{P}\right\}\right\rangle&\left\langle(\Delta\hat{P})^{2}\right\rangle\end{bmatrix}$, (62)
@@ Appendix B The Helstrom bound and SQL of S-BPSK under phase diffusion condition
$P_{\text{SQL}}^{\text{S-BPSK}}\left(N,\sigma\right)$ $=\int_{0}^{+\infty}\mathrm{d}x\cdot P\left(x|\rho_{i}^{\text{pd}}\right)$ (65)
@@
$=\int_{\mathbf{R}}\mathrm{d}\phi\cdot g_{\sigma}(\phi)\cdot P_{\text{SQL}}^{\text{S-BPSK}}(\phi)$
@@
$P_{\text{SQL}}^{\text{S-BPSK}}(\phi)=\int_{0}^{+\infty}\mathrm{d}x\cdot P\left(x|\rho_{i,\phi}\right)=\int_{0}^{+\infty}\mathrm{d}x\int_{\mathbf{R}}\mathrm{d}p\cdot W\left(x,p|\rho_{i,\phi}\right)$ (66)
@@
$P_{\text{HB}}^{\text{S-BPSK}}\left(N,\sigma\right)=\frac{1}{2}\left(1-\left\|P_{0}\rho_{0}^{\text{pd}}-P_{1}\rho_{1}^{\text{pd}}\right\|_{1}\right)$ (67)
where $\left\|A\right\|_{1}=\mathbf{Tr}\sqrt{A^{\dagger}A}$ denotes the trace norm, and $P_{0},P_{1}$ denote prior probabilities. This expression applies to the calculation of the Helstrom bound for binary states, whether pure or mixed.
@@ Acknowledgements.
This work was supported by Independent Innovation Science Fund Program of National University of Defense Technology, China (Grant No. 22-ZZCX-036), Key research&development program of Guangxi (Grant No. GuiKeAB23075155), Innovative Talent Development Fund of Information Support Force Engineering University (Grant No. XJKT-QT-25-02-GW), and the Key Research & Development Program of Guangxi (Grant No. AB23075112).
@@
author contributions
E.B. conceived the study, performed the theoretical analysis and numerical simulations, and wrote the initial manuscript draft. J.P. and T.W. provided critical guidance on the research direction and methodology. C.D. and Z.Z. supervised the project, interpreted the results, and secured the funding. T.W., J.P., K.W., F.S., C.Z. and Y.L. offered expertise in specialized theoretical aspects and reviewed the manuscript.
competing interests
No potential conflict of interest was reported by the authors